Ask
your own question, for FREE!
Mathematics
11 Online
a box contain 9 novels, 7 biographies, and 5 reference books. Alyssa randomly chooses a book, then she randomly chooses another book without replacing the first book what is the probability that she chooses two biographies ? write as decimal
Still Need Help?
Join the QuestionCove community and study together with friends!
P(choosing a biography) = 7/21, since there are 21 books in total and 7 of them are biographies. Once you selected a biography, there are only 6 left, and 20 books total, so the probability of choosing a second biography is 6/20 Hence, P(choosing biography one time AND choosing a biography a second time) = P(choosing biography one time) * P(choosing biography second time) = 7/21 * 6/20 = 1/10 = 0.1
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
5 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal